Grade 12Algebra

In a quiz, team A scored – 40, 10, 0 and team B scored 10, 0, – 40 in three successive rounds. Which team scored more? Can we say that we can add integers in any order?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that
Score of team A = -40, 10, 0
Total score obtained by team A = - 40 + 10 + 0
= - 30
Score of team B = 10, 0, -40
Total score obtained by team B = 10 + 0 + (-40) = 10 + 0–40
= - 30
Thus, the score of the both A team and B team is same.
Yes, we can say that we can add integers in any order

Thanks